40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 29 03:16:21 2016, ericq, Update, LSC, DRFPMI Locked Once More 2016-03-29_calibdarm.pdf2016-03-29_DRFPMI_errctrl.pdfDRFPMI_DTT.zip
    Reply  Wed Mar 30 16:40:24 2016, ericq, Update, LSC, 2016 vs 2010 2016_vs_2010.png
    Reply  Mon Apr 11 16:06:30 2016, ericq, Update, LSC, DRFPMI Data Archived 
Message ID: 12069     Entry time: Mon Apr 11 16:06:30 2016     In reply to: 12053
Author: ericq 
Type: Update 
Category: LSC 
Subject: DRFPMI Data Archived 

I have copied over the complete frame files from two DRFPMI lock acquisitions + locks to /frames/archive. The data should be safe from the wiper script here.

One, under the subfolder DRFPMI_Mar29_cal is the lock where the CAL-DARM channel is properly calibrated at GPS time 1143274087.

The other lock, under DRFPMI_MAR29_nocal, does not have the calibration set up yet, but was a much quicker acquistion (<2 min from ALS acquisition to DRFPMI) and longer lock (~8min).

ELOG V3.1.3-